FILLING-INVARIANTS AT INFINITY FOR

MANIFOLDS OF NONPOSITIVE CURVATURE

Noel Brady Benson Farb

Abstract. In this paper we construct and study isoperimetric functions at infinity for Hadamard manifolds. These quasi-isometry invariants give a measure of the spread of geodesics in such a manifold.

Typeset by AMS-TEX